A tin surface emits electrons after shining with a light of frequency of 2.61×1015 s-1. Calculate the maximum kinetic energy of the ejected electron if the work function of tin is 4.09 eV.
Answer #1

Given:

Eo =4.09 eV

=4.09*1.602*10^-19 J

=6.552*10^-19 J

Given:

f = 2.61*10^15 s-1

use:

E = h*f

=(6.626*10^-34 J.s)*(2.61*10^15) s-1

= 1.729*10^-18 J

use:

E = Eo + kinetic energy

1.729*10^-18 = 6.552*10^-19 + kinetic energy

kinetic energy = 1.074*10^-18 J

Answer: 1.07*10^-18 J